You are a manager at Glass Inc. a mirror and window supplier. Recently, you conducted

a study of the production process for your single-side encapsulated window. The results

from the study are summarized below and are based on the eight units of capital

currently available at your plant. Workers are paid $60 per unit, per unit capital costs

are $20, and your encapsulated windows sell for $12 each. Given this information,

optimize your human resource and production decisions. Do you anticipate earning a

profit or loss? Explain carefully.

Labor Output

0 0

1 10

2 30

3 60

4 80

5 90

6 95

7 95

8 90

9 80

10 60

11 30


1
Expert's answer
2022-04-15T12:37:59-0400

There'll are currently 8 units of capital available at the plant.


labour output MP AP

0 0 0 0

1 10 10-0=10 10/1=10

2 30 30-10=20 30/2=15

3 60 60-30=30 60/3=20

4 80 80-60=20 80/4=20

5 90 90-80=10 90/5=14

6 95 95-90=5 95/6=15.8

7 95 95-95=0 95/7=13.5

8 90 90-95=-5 90/8=11.25

9 80 80-90=-10 80/9=8.89

10 60 60-80=-20 60/10=6

11 30 30-60=-30 30/11=2.9


Maximizing productivity on 3, 4 workers

Profit of maximization:

"80 \cdot 12-60\cdot4-20 \cdot 8 = 560"